OpenStudy (anonymous):

HELP PLEASE!! Solve for y (y+1)^2=1

OpenStudy (anonymous):

(y+1)^2 =1 sqrt((y+1)^2) =sqrt1 y+1=1 y=0

OpenStudy (anonymous):

Oh yea zero my bad :/

OpenStudy (anonymous):

I tried, webassign said that's wrong(

OpenStudy (anonymous):

The other root is y + 1 = -1 -> y = -2.